OpenStudy (anonymous):

Solve the equation x-6/2 = 3 for x.

OpenStudy (anonymous):

x-6=6 x=12

OpenStudy (anonymous):

x-6 = 3 *2=6 x-6=6 x=12

mathslover (mathslover):

Let the user communicate @sai1234 .. This is a site for live to live solutions

OpenStudy (lgbasallote):

\[\frac{x-6}{2} = 3\] multiply both sides by 2 \[\frac{x-6}{2} \times 2 = 3 \times 2\] \[x - 6 = 3 \times 2\] \[x- 6 = 6\] \[x -6 + 6 = 6+6\] \[x = 6+ 6\] \[x = 12\]

mathslover (mathslover):

Please let the user do the questions himself / herself ..

mathslover (mathslover):

or post such answers like @lgbasallote posted..

OpenStudy (anonymous):

x=12

OpenStudy (anonymous):

@mathslover OK. I'm definitely trying. I'm only asking questions that I'm completely stuck on. I have about 4o questions that I'm doing with most of them completed except a few

mathslover (mathslover):

Hmn @mimosa101 No problem ... but may i know that ..r u getting some common problems if yes then have a look to this : http://openstudy.com/study#/updates/4fe0b1c0e4b06e92b86fd82c

OpenStudy (anonymous):

Thanks for everyone help